OpenStudy (anonymous):

I have a math question that reads: y = -3x2 - 6x +2 , find and graph. I have to find several things like the vertex , y-intercept, and the x intercept. However what I need to know is how do i find the y intercept and x intercept.

OpenStudy (anonymous):

y intercept is the constant. it is 2

OpenStudy (anonymous):

x intercept is what you get when you solve \[-3x^2-6x+2=0\]

OpenStudy (anonymous):

y intecept is 2 x intercept is factoring -3x2-6x+2 vertex is where -6x-6=0 which is -1

OpenStudy (anonymous):

first coordinate of the vertex is \[-\frac{b}{2a}=-\frac{-6}{2\times -3}=-1\]

OpenStudy (anonymous):

-6x-6=0 was found using the derivative which you will learn in calculus though

OpenStudy (anonymous):

right, but you always use \[-\frac{b}{2a}\] anyway you are going to have to solve \[-3x^2-6x+2=0\] with the quadratic formula because it does not factor over the integers

OpenStudy (anonymous):

you will get \[\frac{-3\pm\sqrt{15}}{3}\]

OpenStudy (anonymous):

Thank you! I got it!